The floor of a shed given on the right has an area of 75 square feet. The floor is in the shape of a rectangle whose length is 5 feet less than twice the width. Find the length and the width of the floor of the shed.

Answer by ikleyn(52779)

.

            MENTAL   SOLUTION


Let the width be w feet; then the length is  (2w-5) feet.


The equation for the area is


    w*(2w-5) = 75.


Double both sides


    2w*(2w-5) = 150.


Two numbers that give the product 150 and differ in 5 units are 10 and 15.


So, 2w = 15;  hence, w = 7.5.


ANSWER.  The width is 7.5 ft;  the length is 2*7.5-5 = 10 ft.

Solved.